OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single?

OpenStudy (anonymous):

Ok put like terms together and solve for x first. \[x-0.6x=384\] The value of x will be the total number of people who downloaded songs so find 60% of x.

OpenStudy (anonymous):

so abscially 0.6x =384?

OpenStudy (anonymous):

No. Solve for x first. it would be x= 384/0.4

OpenStudy (anonymous):

Then take 60 percent of that.

OpenStudy (anonymous):

Solve for x first. I put the like terms together for you. Now solve for x. Then find 60% of x. Which is \[0.6 \times x\]

OpenStudy (anonymous):

umm what?

OpenStudy (anonymous):

Look at my first comment carefully. Solve for x in the equation.

OpenStudy (anonymous):

Then multiply the value you get for x by 60%

OpenStudy (anonymous):

Dude. Solve x first. When you simpify the original equation you get 0.4x=384. Divide each side by 0.4. Then whatever you get x for, multiply it by 0.6 and there you have your answer

OpenStudy (anonymous):

576?

OpenStudy (anonymous):

Is that a question?

OpenStudy (anonymous):

my answer

OpenStudy (anonymous):

Thats correct